# On the nullspace of split graphs

### v1

@@ 6 Split graphs with nullity 1
    The equality $\operatorname{Supp}(Sp^{\prime})\cap K=\operatorname{Supp}(Sp)\cap K$ can easily be derived from Lemma 6.8. To obtain that $\operatorname{Supp}(Sp^{\prime})\cap S^{\prime}\subset\operatorname{Supp}(Sp)\cap S^{\prime}$, recall that, by item (2) of Theorem 3.4 and item (1) of Theorem 3.2, there exists a (unique) vector $y_{0}$ such that $(I-J)^{-1}Py_{0}=z$. Then, $\operatorname{Supp}(Sp^{\prime})\cap S^{\prime}=\operatorname{Supp}(y_{0})$. On the other hand, we have $(I-J)^{-1}R(y_{0},0)^{t}=z$ as well, which means that $(z,y_{0},0)^{t}\in\operatorname{\mathcal{N}}(Sp)$. Hence, $\operatorname{Supp}(y_{0})\subset\operatorname{Supp}(Sp)\cap S^{\prime}$.
@@
    Obviously, $\langle B_{s}\rangle\subset\operatorname{\mathcal{R}}(A(Sp))$. Since $s\in\operatorname{Supp}(Sp)$ by (1), there exists $x\in\operatorname{\mathcal{N}}(Sp)$ such that $x_{s}\neq 0$ and $x_{v}=0$ for all $v\in V(Sp)\setminus V(Sp^{\prime})$ (recall the proof of (1)). Therefore, we can solve for $A(Sp)_{*s}$ from the equality
    $0=\sum_{j\in V(Sp)}x_{j}A(Sp)_{*j}=x_{s}A(Sp)_{*s}+\sum_{j\in V(Sp^{\prime})\setminus\{s\}}x_{j}A(Sp)_{*j},$
    and show that $A(Sp)_{*s}\in\langle B_{s}\rangle$. For each $v\in V(Sp)\setminus V(Sp^{\prime})$, it is clear by definition of $P$ that $A(Sp)_{*v}$ can be written as a nontrivial linear combination of elements in $\{A(Sp)_{*j}:j\in S^{\prime}\}$ (if $s\in S^{\prime}$, recall that $A(Sp)_{*s}\in\langle B_{s}\rangle$). Hence, $\langle B_{s}\rangle=\operatorname{\mathcal{R}}(A(Sp))$.
@@
    By (2), $\operatorname{rk}(A(Sp))=|B_{s}|=|V(Sp^{\prime})|-1$. Moreover, $|V(Sp^{\prime})|=|S^{\prime}|+|K|$ and $|S^{\prime}|=\operatorname{rk}(R)$, by the definition of $Sp^{\prime}$.
@@
    $0=P(-y_{v})+1R_{*v}+\sum_{j\in W\setminus\{v\}}0R_{*j}=R(-y_{v},e_{v})^{t},$
    we see that $(-y_{v},e_{v})\in\operatorname{\mathcal{N}}(R)$. Note that $\{(-y_{v},e_{v}):v\in W\}$ is clearly a basis for $\operatorname{\mathcal{N}}(R)$ (recall that $|W|=\operatorname{nl}(R)$). Consequently, we get the desired result by repeating the same argument used in the proof of Theorem 5.3.
Clearly, the vector $z$ in Theorem 6.9 can always be chosen with integer coordinates such that $\gcd\{z_{i}:i\in\operatorname{Supp}(z)\}=1$. In such a case, observe that $z$ becomes, up to multiplication by 1-1, an interesting algebraic invariant of split graphs with supported clique vertices.
@@ 7 Tyshkevich composition
The behavior of the nullspace under graph operations is a natural and important line of inquiry. In this section, we analyze how the nullspace of a split graph interacts with the Tyshkevich composition, a well-known product that builds a larger graph by fully joining the clique of a split graph to an arbitrary second graph. We establish fundamental results showing how the singularity of the composition is controlled by the properties of its components, providing a powerful tool for constructing graphs with prescribed nullity and for understanding the propagation of kernel vectors through this structured gluing.
@@
This operation was introduced by R. Tyshkevich in [tyshkevich2000decomposition]. In Figure 6 we can see an example of this composition.
@@
SSGG
Figure 6: Tyshkevich’s composition of $S\approx P_{4}$ and $G\approx C_{4}$.
The Tyshkevich composition preserves (and potentially amplifies) the singularity coming from the bipartite part $R$ of the split graph.
@@
In particular, if $\operatorname{\mathcal{N}}(R)\neq 0$, then $Sp\circ G$ is singular.
@@
Since $(K\cup K^{\prime},S\cup S^{\prime})$ is a bipartition for $H=Sp\circ Sp^{\prime}$, we can label $V(H)$ following the ordering $(K,K^{\prime},S,S^{\prime})$, and write
@@ 8 Determinant of split graphs
For split graphs, the block structure of the adjacency matrix allows us to derive explicit combinatorial formulas for the determinant, using Schur complement and the matrix determinant lemma.
In this section, we establish a precise formula for $\det(Sp)$ in terms of the bipartite adjacency matrix $R$ and its row sums, providing a direct computational method to determine the singularity of a split graph and revealing how the determinant is influenced by the interplay between the clique and the independent set.
@@
Let $Sp(K,S)$ be a split graph with $|K|=\omega\geq 2$. If
@@
$\det(Sp)=(-1)^{\omega-1}(\omega-1)\det\left(R^{t}R-\frac{rr^{t}}{\omega-1}\right).$
Furthermore, if $\operatorname{\mathcal{N}}(R)=0$, we have
$\det\left(R^{t}R-\frac{rr^{t}}{\omega-1}\right)=\left(1-\frac{1}{\omega-1}r^{t}(R^{t}R)^{-1}r\right)\det(R^{t}R).$
@@
We apply the Schur complement (Theorem 8.1) with A=JIA=J-I, $B=R$, $C=R^{t}$, and $D=0$:
$\det(A(Sp))=\det(J-I)\det(0-R^{t}(J-I)^{-1}R)=(-1)^{\omega-1}(\omega-1)\det(-R^{t}(J-I)^{-1}R).$
Using the fact that $(I-J)^{-1}=I-(\omega-1)^{-1}J$, we have:
$R^{t}(J-I)^{-1}R=-R^{t}(I-J)^{-1}R=-R^{t}\left(I-\frac{1}{\omega-1}J\right)R=-R^{t}R+\frac{1}{\omega-1}R^{t}JR.$
@@
$\det(A(Sp))=(-1)^{\omega-1}(\omega-1)\det\left(R^{t}R-\frac{rr^{t}}{\omega-1}\right).$
For the second part, if $\operatorname{\mathcal{N}}(R)=0$, then $R^{t}R$ is invertible, and we apply the matrix determinant lemma (Theorem 8.2) with $M=R^{t}R$ and $u=v=-\frac{1}{\omega-1}r$:
$\det\left(R^{t}R-\frac{rr^{t}}{\omega-1}\right)=\left(1-\frac{1}{\omega-1}r^{t}(R^{t}R)^{-1}r\right)\det(R^{t}R).\qed$
@@
Let $Sp(K,S)$ be a split graph such that $|K|\geq 2$ and $\operatorname{\mathcal{N}}(R)=0$. Then, $Sp$ is singular if and only if
@@
From Theorem 8.3, SpSp is singular if and only if
$1-\frac{1}{|K|-1}r^{t}(R^{t}R)^{-1}r=0.$
9 The intersection graph
This section introduces the intersection graph of an split graph. It captures the overlap pattern of neighborhoods in the independent set, and its connectivity is equivalent to the connectivity of the bipartite part of the split graph.
Let $\mathcal{N}=\{N_{v}:v\in S\}$ be a finite multiset of finite non-empty sets such that $\left(\bigcup_{v\in S}N_{v}\right)\cap S=\varnothing$ and $N_{v}\neq\bigcup_{u\in S}N_{u}$ for all $v\in S$. Note that $\mathcal{N}$ induces the balanced split graph $(Sp_{\mathcal{N}},K,S)$, where $K=\bigcup_{v\in S}N_{v}$ and $N_{v}=N_{Sp}(v)$, for each $v\in S$. Conversely, to every balanced split graph $(Sp,K^{\prime},S^{\prime})$, we can associate the multiset $\mathcal{N}_{Sp}=\{N_{Sp}(v):v\in S^{\prime}\}$. The intersection graph of a balanced split graph $Sp(K,S)$, denoted by $\operatorname{Int}(Sp)$, is the (simple) graph with $S$ as vertex set, where $uv\in E(\operatorname{Int}(Sp))$ if and only if $N_{Sp}(u)\cap N_{Sp}(v)\neq\varnothing$.
If $Sp(K,S)$ is a split graph and $E(K)=\{xy\in E(Sp):x,y\in K\}$, then SpE(K)Sp-E(K) is a bipartite graph with bipartition $(K,S)$.
Proposition 9.1.
For every balanced split graph $Sp(K,S)$, $\operatorname{Int}(Sp)$ is connected if and only if SpE(K)Sp-E(K) is connected.
Proof.
Assume that $\operatorname{Int}(Sp)$ is connected and take two distinct vertices $x,y\in B=Sp-E(K)$. We have to analyze 3 cases:
  1. 1.
    $x,y\in S$: There is a path in $\operatorname{Int}(Sp)$ connecting $x$ to $y$. For each edge $uv$ in this path, there exists $k\in K$ such that $uk,kv\in E(B)$, giving a path in $B$ from $x$ to $y$.
  2. 2.
    $x\in K,y\in S$: Since $x$ has a neighbor $x^{\prime}\in S$ and $\operatorname{Int}(Sp)$ is connected, there is a path from $x^{\prime}$ to $y$ in $\operatorname{Int}(Sp)$, which lifts to a path in $B$ from $x$ to $y$.
  3. 3.
    $x,y\in K$: Each has neighbors in $S$, and by case (1) those neighbors are connected in $B$, giving a path from $x$ to $y$.
Conversely, suppose that $B$ is connected and take two distinct vertices $x,y\in\operatorname{Int}(Sp)$. By assumption, there is a path $P\subset B$ from $x$ to $y$. Since $B$ is bipartite, $P$ alternates vertices of $S$ and $K$, i.e., $P=xa_{1}b_{2}\ldots b_{r-1}a_{r}y$, where $\{x,b_{2},b_{4},\ldots,b_{r-1},y\}\subset S$ and $\{a_{1},a_{3},\ldots,a_{r}\}\subset K$. Consequently, $xb_{2}b_{4}\ldots b_{r-1}y$ is a path connecting $x$ to $y$ in $\operatorname{Int}(Sp)$. ∎
Proposition 9.2.
For every balanced split graph $Sp(K,S)$,
$\max\{|N_{Sp}(x)\cap S|:x\in K\}\leq\omega\big(\operatorname{Int}(Sp)\big).$
Proof.
If $x\in K$ and $u,v\in N_{Sp}(x)\cap S$, then $x\in N_{Sp}(u)\cap N_{Sp}(v)$, which means that $uv\in E(\operatorname{Int}(Sp))$. Hence, $N_{Sp}(x)\cap S$ is a clique in $\operatorname{Int}(Sp)$. ∎
If $G$ is a nontrivial graph, recall that $G$ is triangle-free if and only if $\omega(G)=2$.
Lemma 9.3.
Suppose that $\omega(G)=2$, where $G=\operatorname{Int}(S)$. If $x\in\operatorname{\mathcal{N}}(R)$ and $x_{v}=0$ for some $v\in V(G)$, then $x_{u}=0$ for all $u\in N_{G}(v)$.
Proof.
If $uv\in E(G)$, consider a vertex $w\in N_{S}(u)\cap N_{S}(v)$. Since $0=R_{w*}x=x_{u}+x_{v}$ and $x_{v}=0$, then $x_{u}=0$. ∎
We finish the paper with proving a complete classification of balanced split graphs with connected, triangle-free intersection graphs. It shows that in this case, the nullity is at most 1, and we have a clear dichotomy based on whether some clique vertex has only one neighbor in the independent set. This structural result connects combinatorial properties (neighborhood sizes, intersection graph structure) with spectral properties (nullity).
Theorem 9.4.
Let $Sp(K,S)$ be a balanced split graph. Suppose that $\operatorname{Int}(Sp)$ is connected and $\omega\big(\operatorname{Int}(Sp)\big)=2$. Then, we have the following:
  1. 1.
    if there is a $v\in K$ such that $|N_{Sp}(v)\cap S|=1$, then $\operatorname{\mathcal{N}}(R)=0$ and $\operatorname{nl}(Sp)\leq 1$;
  2. 2.
    otherwise, we have that $|N_{Sp}(v)\cap S|=2$ for all $v\in K$, $\operatorname{Supp}(Sp)\subset S$ and $\operatorname{nl}(Sp)\leq 1$.
Proof.
Let $G=\operatorname{Int}(Sp)$. Since $G$ is connected and $\omega(G)=2$, we know from Proposition 9.2 and Proposition 9.1 that $|N_{Sp}(v)\cap S|\in\{1,2\}$, for all $v\in K$.
If $x\in\operatorname{\mathcal{N}}(R)$ and $x_{v}=0$ for some $v\in S$, then $x=0$, by Lemma 9.3 and by connectedness of $G$. In particular, this means the following:
  1. 1.
    If $|N_{Sp}(v)\cap S|=1$ for some $v\in K$, then $\operatorname{\mathcal{N}}(R)=0$ (i.e., if $\operatorname{\mathcal{N}}(R)\neq 0$, then $|N_{Sp}(v)\cap S|=2$ for all $v\in K$);
  2. 2.
    If $x\in\operatorname{\mathcal{N}}(R)\setminus\{0\}$, then $x_{v}\neq 0$ for all $v\in S$.
When $\operatorname{\mathcal{N}}(R)=0$, $Sp$ is nonsingular if and only if $\operatorname{Cker}{\left(Sp\right)}=0$, by item (2) of Theorem 3.2 and item (3) of Theorem 3.4. Then, $\operatorname{nl}(Sp)\leq 1$, by Theorems 5.2 and 5.3.
When $|N_{Sp}(v)\cap S|=2$ for all $v\in K$, we know from item (1) of Theorem 3.6 that $\operatorname{Supp}(Sp)\subset S$. So, $\operatorname{nl}(Sp)=\operatorname{nl}(R)$, by item (3) of Theorem 3.2.
Assume that $\operatorname{nl}(R)\geq 2$ and let $x,y$ be two independent vectors in $\operatorname{\mathcal{N}}(R)\setminus\{0\}$. Recall that every entry of $x$ and $y$ is nonzero, by (2). If $a,b$ are the first entries of $x,y$ respectively, then the first entry of
$z=a^{-1}x-b^{-1}y\in\operatorname{\mathcal{N}}(R)$
is zero. Hence, $z=0$ by (2), which contradicts the assumption that $x,y$ are linearly independent. Therefore, $\operatorname{nl}(R)\leq 1$. ∎
@@ Acknowledgments
This work was partially supported by Universidad Nacional de San Luis (Argentina), grant PROICO 03-0723, Agencia I+D+i, grants PICT-2020-Serie A-00549 and PICT-2021-CAT-II-00105, CONICET grant PIP 11220220100068CO.

### v2

@@ 6 Split graphs with nullity 1
    The equality $\operatorname{Supp}(Sp^{\prime})\cap K=\operatorname{Supp}(Sp)\cap K$ can easily be derived from Lemma 6.8. To obtain that $\operatorname{Supp}(Sp^{\prime})\cap S^{\prime}\subset\operatorname{Supp}(Sp)\cap S^{\prime}$, recall that, by item (2) of Theorem 3.4 and item (1) of Theorem 3.2, there exists a (unique) vector $y_{0}$ such that $(I-J)^{-1}Py_{0}=z$. Then,
    $\operatorname{Supp}(Sp^{\prime})\cap S^{\prime}=\{u\in S^{\prime}:(y_{0})_{u}\neq 0\}.$
    On the other hand, we have $(I-J)^{-1}R(y_{0},0)^{t}=z$ as well, which means that $(z,y_{0},0)^{t}\in\operatorname{\mathcal{N}}(Sp)$. Hence,
    $\{u\in S^{\prime}:(y_{0})_{u}\neq 0\}\subset\operatorname{Supp}(Sp)\cap S^{\prime}.$
@@
    Obviously, $\langle B_{s}\rangle\subset\operatorname{\mathcal{R}}(Sp)$. Since $s\in\operatorname{Supp}(Sp)$ by (1), there exists $x\in\operatorname{\mathcal{N}}(Sp)$ such that $x_{s}\neq 0$ and $x_{v}=0$ for all $v\in V(Sp)\setminus V(Sp^{\prime})$ (recall the proof of (1)). Therefore, we can solve for $A(Sp)_{*s}$ from the equality
    $0=\sum_{j\in V(Sp)}x_{j}A(Sp)_{*j}=x_{s}A(Sp)_{*s}+\sum_{j\in V(Sp^{\prime})\setminus s}x_{j}A(Sp)_{*j},$
    and show that $A(Sp)_{*s}\in\langle B_{s}\rangle$. For each $v\in V(Sp)\setminus V(Sp^{\prime})$, it is clear by definition of $P$ that $A(Sp)_{*v}$ can be written as a nontrivial linear combination of elements in $\{A(Sp)_{*j}:j\in S^{\prime}\}$ (if $s\in S^{\prime}$, recall that $A(Sp)_{*s}\in\langle B_{s}\rangle$). Hence, $\langle B_{s}\rangle=\operatorname{\mathcal{R}}(Sp)$.
@@
    By (2), $\operatorname{rk}(Sp)=|B_{s}|=|V(Sp^{\prime})|-1$. Moreover, $|V(Sp^{\prime})|=|S^{\prime}|+|K|$ and $|S^{\prime}|=\operatorname{rk}(R)$, by the definition of $Sp^{\prime}$.
@@
    $0=P(-y_{v})+1R_{*v}+\sum_{j\in W\setminus v}0R_{*j}=R(-y_{v},e_{v})^{t},$
    we see that $(-y_{v},e_{v})^{t}\in\operatorname{\mathcal{N}}(R)$. Note that $\{(-y_{v},e_{v}):v\in W\}$ is clearly a basis for $\operatorname{\mathcal{N}}(R)$ (recall that $|W|=\operatorname{nl}(R)$). Consequently, we get the desired result by repeating the same argument used in the proof of Theorem 5.3. ∎
Clearly, the vector $z$ constructed in Theorem 6.9 can always be chosen with integer coordinates such that $\gcd\{z_{i}:z_{i}\neq 0\}=1$. In such a case, observe that $z$ becomes, up to multiplication by 1-1, an interesting algebraic invariant of split graphs with supported clique vertices.
@@ 7 Tyshkevich composition
Understanding how the nullspace behaves under graph operations is a natural and important question. In this section, we analyze how the nullspace of a split graph interacts with the Tyshkevich composition, a well-known product that builds a larger graph by fully joining the clique of a split graph to an arbitrary second graph. We establish fundamental results showing how the singularity of the composition is controlled by the properties of its components, providing a powerful tool for constructing graphs with prescribed nullity and for understanding the propagation of kernel vectors through this structured gluing.
@@
This operation was introduced by R. Tyshkevich in [19]. In Figure 6 we can see an example of this composition.
@@
SpSpGG
Figure 6: Tyshkevich composition of $Sp\approx P_{4}$ and $G\approx C_{4}$.
The Tyshkevich composition preserves (and potentially amplifies) the singularity coming from the biadjacency matrix $R$ of the split graph.
@@
In particular, if $\operatorname{\mathcal{N}}(R)\neq\{0\}$, then $Sp\circ G$ is singular.
@@
Since $(K\cup K^{\prime},S\cup S^{\prime})$ is an s-partition for $H=Sp\circ Sp^{\prime}$, we can label $V(H)$ following the ordering $(K,K^{\prime},S,S^{\prime})$, and write
@@ 8 Determinant of split graphs
Thanks to the block structure of their adjacency matrices, split graphs admit explicit combinatorial determinant formulas that can be obtained with the Schur complement and the matrix-determinant lemma.
In this section, we establish a precise formula for $\det(Sp)$ (i.e., $\det(A(Sp))$) in terms of the biadjacency matrix $R$, providing a direct computational method to determine the singularity of a split graph and revealing how the determinant is influenced by the interplay between the clique and the independent set.
@@
Let $Sp(K,S)$ be a split graph with $|K|=k\geq 2$. If
@@
$\det(Sp)=(-1)^{k-1}(k-1)\det\left(R^{t}R-\frac{rr^{t}}{k-1}\right).$
Furthermore, if $\operatorname{\mathcal{N}}(R)=\{0\}$, we have
$\det\left(R^{t}R-\frac{rr^{t}}{k-1}\right)=\left(1-\frac{1}{k-1}r^{t}(R^{t}R)^{-1}r\right)\det(R^{t}R).$
@@
We apply the Schur complement (Theorem 8.1) with
A=JIA=J-I, $B=R$, $C=R^{t}$, and $D=0$:
$\det(Sp)=\det(J-I)\det(0-R^{t}(J-I)^{-1}R)=$
$(-1)^{k-1}(k-1)\det(-R^{t}(J-I)^{-1}R).$
Using the fact that $(I-J)^{-1}=I-(k-1)^{-1}J$, we have:
$R^{t}(J-I)^{-1}R=-R^{t}(I-J)^{-1}R=-R^{t}\left(I-\frac{1}{k-1}J\right)R=$
$R^{t}R+\frac{1}{k-1}R^{t}JR.$
@@
$\det(Sp)=(-1)^{k-1}(k-1)\det\left(R^{t}R-\frac{rr^{t}}{k-1}\right).$
For the second part, if $\operatorname{\mathcal{N}}(R)=\{0\}$, then $R^{t}R$ is invertible, and we apply the matrix determinant lemma (Theorem 8.2) with $M=R^{t}R,u=\frac{r}{1-k}$ and $v=r$.
$\det\left(R^{t}R-\frac{rr^{t}}{k-1}\right)=\left(1-\frac{1}{k-1}r^{t}(R^{t}R)^{-1}r\right)\det(R^{t}R).\qed$
@@
Let $Sp(K,S)$ be a split graph such that $|K|\geq 2$ and $\operatorname{\mathcal{N}}(R)=\{0\}$. Then, $Sp$ is singular if and only if
@@
It immediately follows from Theorem 8.3, because $\det(R^{t}R)\neq 0$. ∎
@@ Acknowledgments
This work was partially supported by Universidad Nacional de San Luis (Argentina), grant PROICO 03-0723, Agencia I+D+i, grant PICT-2020-Serie A-00549, CONICET, grant PIP 11220220100068CO.